OpenStudy (anonymous):

A data set consists of the following data points: (3,5), (5,8), (6,13) The slope of the best fit line is 2.5. Find the y-intercept of this line.

OpenStudy (anonymous):

find xbar and ybar. the line goes through that point and has the given slope. from there you can write the point-slope form of the line and solve for y to get it into slope-intercept form and thus find the intercept

OpenStudy (anonymous):

whats xbar ? and y \bar?

OpenStudy (anonymous):

ybar

OpenStudy (anonymous):

the averages of the x's & y's, respectively

OpenStudy (anonymous):

okay... but how do i found the y intercept of 2.5?

OpenStudy (anonymous):

xbar\[=\frac{ 3+5+6 }{ 3 }\]

OpenStudy (anonymous):

do you remember linear equations? y = mx+b

OpenStudy (anonymous):

sort of,

OpenStudy (anonymous):

4.6 ?

OpenStudy (anonymous):

so m is the slope and b is the y-intercept. the same will hold for regression lines except the data is used to get the slope and y-intercept

OpenStudy (anonymous):

so that's xbar, find ybar

OpenStudy (anonymous):

8.6

OpenStudy (anonymous):

now you have the point (4.666666, 8.666666) and a slope of 2.5 the point-slope form of a line is \[\left( y-y _{1} \right)=m \left( x-x _{1} \right)\] where m is the slope (2.5 in this case) and the point is \[\left( x _{1}, y _{1} \right)\]

OpenStudy (anonymous):

and whats the other x and y ?

OpenStudy (anonymous):

put the numbers in their proper places and then solve for y. the equation will be in the slope-intercept form, i.e.,\[y=mx+b\]and whatever b is will be the intercept you seek

OpenStudy (anonymous):

they are the variables and they stay

OpenStudy (anonymous):

i dont get it,

OpenStudy (anonymous):

how many x and y 's are there?

OpenStudy (anonymous):

you should review linear equations... have a look at this site: http://www.mathsisfun.com/algebra/linear-equations.html

OpenStudy (anonymous):

i dont have time for reviewing, :/ i need to pass this by tomorowwww. well thanks for helping anyway.

OpenStudy (anonymous):

y−8.6=2.5(x−4.6) solve for y and the y intercept will be the constant term (the one not multiplied by x)

OpenStudy (anonymous):

so whats the answer???-.-